# The Obstacle Problem Arising from the American Chooser Option

### v2

@@ 4.2. Properties of the free boundary
Similar results and approaches used in the proof can be found in previous research; see, e.g., [27, 28]. Therefore, we shall omit the details of approximations. We first prove Part (i). Let $\tilde{C}^{A}_{n,\varepsilon}$ and $\tilde{P}^{A}_{n,\varepsilon}$ be the unique solutions of the following equations with Dirichlet boundary conditions:
@@
where $\beta_{c,\varepsilon}$ and $\beta_{p,\varepsilon}$ are appropriate penalty functions, ${\Omega}^{n}_{T_{c}}:=(0,T_{c})\times(-n,n)$ and ${\Omega}^{n}_{T_{p}}:=(0,T_{p})\times(-n,n)$. The existence and uniqueness of the solutions to (4.7) and (4.8) is guaranteed by the Schauder’s fixed-point theorem. Then $\tilde{C}^{A}_{n}:=\lim_{\varepsilon\to 0^{+}}\tilde{C}^{A}_{n,\varepsilon}$ and $\tilde{P}^{A}_{n}:=\lim_{\varepsilon\to 0^{+}}\tilde{P}^{A}_{n,\varepsilon}$ are well-defined and solutions to the following obstacle problems:
@@
since $\beta_{c,\varepsilon}\in C^{\infty}(\mathbb{R})$ is a penalty function satisfying $\beta_{c,\varepsilon}(t)=0$ for all $t\geqslant\varepsilon$ and $e^{x}+2>\varphi_{\varepsilon}(e^{x}-K_{c})+\varepsilon$ holds for all $\varepsilon<1.$ By the comparison principle [6, 52p],
(4.9) $\tilde{C}^{A}_{n,\varepsilon}\leqslant e^{x}+2\ \ \text{for all}\ \ \varepsilon\in(0,1),\ \ \text{hence}\ \ \tilde{C}^{A}_{n}\leqslant e^{x}+2\quad\text{in}\ \ \Omega^{n}_{T_{c}}.$
@@
Again by the comparison principle [6, 52p],
(4.10) $\tilde{P}^{A}_{n,\varepsilon}\leqslant K_{p}+2\ \ \text{for all}\ \ \varepsilon\in(0,1),\ \ \text{hence}\ \ \tilde{P}^{A}_{n}\leqslant K_{p}+2\quad\text{in}\ \ \Omega^{n}_{T_{p}}.$
Moreover, for each $R>0$, ${\lim_{n\to\infty}}\tilde{C}^{A}_{n}$ and ${\lim_{n\to\infty}}\tilde{P}^{A}_{n}$ satisfy (2.5) and (2.6) in the domains $[-R,R]\times(0,T_{c})$ and $[-R,R]\times(0,T_{p})$ respectively. Since the problems (2.5) and (2.6) have unique solutions, taking $n\rightarrow\infty$ in (4.9) and (4.10) yield the results in Part (i).
We next prove Part (ii). Let $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ be the unique solutions to the following equations with Neumann boundary conditions:
@@
Then $\widehat{C}^{A}_{n}:=\lim_{\varepsilon\to 0^{+}}\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n}:=\lim_{\varepsilon\to 0^{+}}\widehat{P}^{A}_{n,\varepsilon}$ are well-defined and solutions to the following obstacle problems:
@@
Note that by $W^{1,2}_{p}$-regularity, $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ are $W^{1,2}_{p}$ functions for all $1<p<\infty$ in the domain $\Omega_{T_{c}}$ and $\Omega_{T_{p}}$, respectively. Thus, the Sobolev embedding theorem implies that $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ are $C^{\frac{\alpha}{2},\alpha}$ functions for some $\alpha\in(0,1)$. Based on the Schauder theory for linear parabolic equations and a bootstrap argument, we deduce that the functions $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ are smooth. Thus, differentiating (4.11) and (4.12) yields that $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}$ and $\partial_{x}\widehat{P}^{A}_{n,\varepsilon}$ satisfies
@@
respectively. Applying the maximum principle for each equation, we deduce $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}\geqslant 0$ in ${\Omega}^{n}_{T_{c}}$ and $\partial_{x}\widehat{P}_{n,\varepsilon}^{A}\leqslant 0$ in ${\Omega}^{n}_{T_{p}}$. Furthermore, since $-\mathcal{L}e^{x}+\beta^{\prime}_{c,\varepsilon}(\cdots)e^{x}\geqslant 0$, we obtain
@@
Therefore, by the maximum principle, we deduce $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}\leqslant e^{x}$ and $\partial_{x}\widehat{P}_{n,\varepsilon}^{A}\geqslant-e^{x}$. Note that $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}$ and $\partial_{x}\widehat{P}^{A}_{n,\varepsilon}$ converge to $\partial_{x}\widehat{C}^{A}_{n}$ and $\partial_{x}\widehat{P}^{A}_{n}$ uniformly in $\Omega_{T_{c}}^{n}$ and $\Omega_{T_{p}}^{n}$, respectively, as $\varepsilon\rightarrow 0^{+}$. Hence $0\leqslant\partial_{x}\widehat{C}^{A}_{n}\leqslant e^{x}$ in $\Omega_{T_{c}}^{n}$ and $-e^{x}\leqslant\partial_{x}\widehat{P}^{A}_{n}\leqslant 0$ in $\Omega_{T_{p}}^{n}$. Finally, passing $n\to\infty$, we obtain the desired inequalities for $\partial_{x}\widehat{C}^{A}$ and $\partial_{x}\widehat{P}^{A}$.
To prove Part (iii), for each fixed $\delta>0$, we first define $\tilde{C}_{n}^{\delta}$ by
@@
Since $\tilde{C}_{n}^{\delta}(0,x)\geqslant(e^{x}-K_{c})^{+}=\tilde{C}^{A}_{n}(0,x)$, by the comparison principle for the variational inequality, (see [7, 80p, problem 5]) we deduce
$\tilde{C}^{A}_{n}(\zeta+\delta,x)=\tilde{C}_{n}^{\delta}(\zeta,x)\geqslant\tilde{C}^{A}_{n}(\zeta,x)$
for all $(\zeta,x)\in\Omega^{n}_{T_{c}-\delta}$. Take $n\rightarrow\infty$ to get
$\tilde{C}^{A}(\zeta+\delta,x)\geqslant\tilde{C}^{A}(\zeta,x)$
for all $(\zeta,x)\in\Omega_{T_{c}-\delta}$. Since $\delta>0$ is arbitrary we conclude $\partial_{\tau}C\geqslant 0$. Similarly, we can obtain $\partial_{\tau}P\geqslant 0$.
@@
Conditions of Schauder fixed point theorem
We show that the function of the operator $\mathcal{F}:\mathcal{D}\to\mathcal{B}$ in the proof of Theorem (3.1) satisfies
@@
    $\mathcal{F}(\mathcal{D})\subset\mathcal{D}$;
@@
    $\mathcal{F}(\mathcal{D})$ is precompact in $\mathcal{B}$.
(1) Fix any $w\in\mathcal{D}$ and consider $u=\mathcal{F}(w).$ Note that $J_{\varepsilon}(\tau,x)\geqslant 0,$ $-\beta_{\varepsilon}(\cdots)\geqslant 0$ and $\ \nu\cdot D_{x}u(\tau,x)\leqslant 0$ for each $(\tau,x)$ at the lateral boundary and the inward pointing normal vector $\nu$ at $(\tau,x).$ Then by the comparison principle [14, 13p], we have $u\geqslant 0.$
@@
[…]
@@
Then the $W^{1,2}_{p}$ estimate combined with the embedding theorem implies that $\|u-u_{j}\|_{C(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}\|\beta_{\varepsilon}^{\prime}(\cdots)(w-w_{j})\|_{L^{\infty}(\Omega_{T}^{n})}$ for some constant $\mathcal{K}>0$. Therefore, it follows that ${\lim_{j\to\infty}u_{j}}=u.$
(3) Let $\{u_{j}\}_{j=1}^{\infty}$ be a sequence in the closure of $\mathcal{F}(\mathcal{D})$ with respect to the $\|\cdot\|_{C(\overline{\Omega_{T}^{n}})}$ norm.
Case 1. For each $u_{j}\in\mathcal{F}(\mathcal{D})$, there exists a sequence of functions $\{w_{j}\}_{j=1}^{\infty}$ such that $\mathcal{F}(w_{j})=u_{j}$. By applying the $W^{1,2}_{p}$ estimate on $u_{j}$, we obtain
$\|u_{j}\|_{W^{1,2}_{p}(\Omega^{n}_{T})}$ $\leqslant\mathcal{K}(\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(w_{j}-J_{\varepsilon})\|_{L^{p}(\Omega^{n}_{T})}+\|e^{n}+e^{-n}\|_{W^{1}_{p}((0,T))})$
$\leqslant\mathcal{K}(\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(-J_{\varepsilon})\|_{L^{p}(\Omega^{n}_{T})}+\|e^{n}+e^{-n}\|_{W^{1}_{p}((0,T))})$
@@
Case 2. For each fixed $u_{i}\notin\mathcal{F}(\mathcal{D})$, there exists a sequence $\{u^{(i)}_{l}\}_{l=1}^{\infty}\subset\mathcal{F}(\mathcal{D})$ such that
@@
Hence, there exists $\{w^{(i)}_{l}\}_{l=1}^{\infty}$ in $\mathcal{D}$ such that $\mathcal{F}(w^{(i)}_{l})=u^{(i)}_{l}$. Again, by the $W^{1,2}_{p}$ estimate, the embedding theorem, and the monotonicity of $\beta_{\varepsilon}$, we deduce $\|u_{l}^{(i)}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$ for some $\alpha\in(0,1)$. Take $l\rightarrow\infty$ to get $\|u_{i}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$.
@@
for all $j\in\mathbb{N}$. By the Arzela-Ascoli theorem there exists a subsequence $\{u_{j_{k}}\}_{k=1}^{\infty}\subset\{u_{j}\}_{j=1}^{\infty}$ and $u\in C(\overline{\Omega_{T}^{n}})$ such that
@@
In particular, $u$ is contained in the closure of $\mathcal{F}(\mathcal{D})$ by the assumption. This proves that $\mathcal{F}(\mathcal{D})$ is precompact in $\mathcal{B}=C(\overline{\Omega_{T}^{n}}).$

### v3

@@ B. Verification
Similar results and approaches used in the proof can be found in previous research; see, e.g., [26, 27]. Therefore, we shall omit the details of approximations.
Proof of Part (i)
Let $\tilde{C}^{A}_{n,\varepsilon}$ and $\tilde{P}^{A}_{n,\varepsilon}$ be a solutions of the following equations with Dirichlet boundary conditions:
@@
where $\beta_{c,\varepsilon}$ and $\beta_{p,\varepsilon}$ are appropriate penalty functions, ${\Omega}^{n}_{T_{c}}:=(0,T_{c})\times(-n,n)$ and ${\Omega}^{n}_{T_{p}}:=(0,T_{p})\times(-n,n)$. It can be seen that the limits $\tilde{C}^{A}_{n}:=\displaystyle\lim_{\varepsilon\to 0^{+}}\tilde{C}^{A}_{n,\varepsilon}$ and $\tilde{P}^{A}_{n}:=\displaystyle\lim_{\varepsilon\to 0^{+}}\tilde{P}^{A}_{n,\varepsilon}$ are well-defined and satisfy the following obstacle problems, respectively:
@@
since $\beta_{c,\varepsilon}\in C^{\infty}(\mathbb{R})$ is a penalty function satisfying $\beta_{c,\varepsilon}(\lambda)=0$ for all $\lambda\geqslant\varepsilon$ and $e^{x}+2>\varphi_{\varepsilon}(e^{x}-K_{c})+\varepsilon$ holds for all $\varepsilon<1.$ Therefore, for each $0<\varepsilon<1$, we deduce by the comparison principle [8, 52p] that $\tilde{C}^{A}_{n,\varepsilon}\leqslant e^{x}+2$ in $\Omega^{n}_{T_{c}}$.
@@
By the comparison principle [8, 52p], $\tilde{P}^{A}_{n,\varepsilon}\leqslant K_{p}+2$ in $\Omega^{n}_{T_{p}}$ holds for all $0<\varepsilon<1$. Take $\varepsilon\to 0^{+}$ to get
$\tilde{C}^{A}_{n}\leqslant e^{x}+2\quad\text{in }\ \Omega_{T_{c}}^{n}\quad\text{and}\quad\tilde{P}^{A}_{n}\leqslant K_{p}+2\quad\text{in }\ \Omega_{T_{p}}^{n}.$
It can be seen further that ${\lim_{n\to\infty}}\tilde{C}^{A}_{n}$ and ${\lim_{n\to\infty}}\tilde{P}^{A}_{n}$ satisfy (2.5) and (2.6) on each compact subset of $\Omega_{T_{c}}$ and $\Omega_{T_{p}}$, respectively. Since the solution of each problem (2.5) and (2.6) is unique, we see that $C^{A}=\displaystyle{\lim_{n\to\infty}}\tilde{C}^{A}_{n}$ in $\Omega_{T_{c}}$ and $P^{A}=\displaystyle{\lim_{n\to\infty}}\tilde{P}^{A}_{n}$ in $\Omega_{T_{p}}$. Therefore, taking $n\rightarrow\infty$ yields the results in Part (i).
Proof of Part (ii)
Let $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ be the unique solutions to the following equations with Neumann boundary conditions:
@@
Then $\widehat{C}^{A}_{n}:=\displaystyle\lim_{\varepsilon\to 0^{+}}\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n}:=\displaystyle\lim_{\varepsilon\to 0^{+}}\widehat{P}^{A}_{n,\varepsilon}$ are well-defined and solutions to the following obstacle problems:
@@
Note that $\widehat{C}^{A}_{n}$ and $\widehat{P}^{A}_{n}$ are obtained by letting $\varepsilon\to 0^{+}$, where $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ converge weakly in $W^{1,2}_{p}$ for each $1<p<\infty$ in the domains $\Omega_{T_{c}}$ and $\Omega_{T_{p}}$, respectively. This yields that $\widehat{C}^{A}_{n}\in W^{1,2}_{p}(\Omega_{T_{c}})$ and $\widehat{P}^{A}_{n}\in W^{1,2}_{p}(\Omega_{T_{p}})$ for each $1<p<\infty$ and the Sobolev embedding theorem further implies that $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ are $C^{\frac{\alpha}{2},\alpha}$ functions for some $\alpha\in(0,1)$. Based on the Schauder theory for linear parabolic equations and a bootstrap argument, we deduce that the functions $\widehat{C}^{A}_{n,\varepsilon}$ and $\widehat{P}^{A}_{n,\varepsilon}$ are smooth. Thus, differentiating (4.16) and (4.17) yields that $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}$ and $\partial_{x}\widehat{P}^{A}_{n,\varepsilon}$ satisfies
@@
respectively. Applying the maximum principle for each equation, we deduce $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}\geqslant 0$ in ${\Omega}^{n}_{T_{c}}$ and $\partial_{x}\widehat{P}_{n,\varepsilon}^{A}\leqslant 0$ in ${\Omega}^{n}_{T_{p}}$. Furthermore, since $\mathcal{L}e^{x}+\beta^{\prime}_{c,\varepsilon}(\cdots)e^{x}\geqslant 0$, we obtain
@@
Therefore, by the maximum principle, we deduce $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}\leqslant e^{x}$ and $\partial_{x}\widehat{P}_{n,\varepsilon}^{A}\geqslant-e^{x}$. Note by the Sobolev embedding theorem [15, Chapter II. Lemma 3.3] that $\partial_{x}\widehat{C}^{A}_{n,\varepsilon}$ and $\partial_{x}\widehat{P}^{A}_{n,\varepsilon}$ converge to $\partial_{x}\widehat{C}^{A}_{n}$ and $\partial_{x}\widehat{P}^{A}_{n}$ uniformly in $\Omega_{T_{c}}^{n}$ and $\Omega_{T_{p}}^{n}$, respectively, as $\varepsilon\rightarrow 0^{+}$. Hence
$0\leqslant\partial_{x}\widehat{C}^{A}_{n}\leqslant e^{x}\quad\text{in }\ \Omega_{T_{c}}^{n}\quad\text{and}\quad-e^{x}\leqslant\partial_{x}\widehat{P}^{A}_{n}\leqslant 0\quad\text{in }\ \Omega_{T_{p}}^{n}.$
Finally, by the uniqueness, passing $n\to\infty$ yields the desired inequalities for $\partial_{x}{C}^{A}$ and $\partial_{x}{P}^{A}$.
Proof of Part (iii)
For each fixed $\delta>0$, we first define $\tilde{C}_{n}^{\delta}$ by
@@
Since $\tilde{C}_{n}^{\delta}(0,x)\geqslant(e^{x}-K_{c})^{+}=\tilde{C}^{A}_{n}(0,x)$, by the comparison principle for the variational inequality, (see [25, Theorem 1.2]) we deduce
$\tilde{C}^{A}_{n}(\zeta+\delta,x)=\tilde{C}_{n}^{\delta}(\zeta,x)\geqslant\tilde{C}^{A}_{n}(\zeta,x)\quad\text{for all }\ (\zeta,x)\in\Omega^{n}_{T_{c}-\delta}.$
Take $n\rightarrow\infty$ and the uniqueness result yields that
${C}^{A}(\zeta+\delta,x)\geqslant{C}^{A}(\zeta,x)\quad\text{for all }\ (\zeta,x)\in\Omega_{T_{c}-\delta}.$
Observe that ${C}^{A}$ is differentiable almost everywhere with respect to $\tau$ in $\Omega_{T_{c}}$, and that $\partial_{\tau}C^{A}\geqslant 0$ almost everywhere. In particular, this condition satisfies the hypothesis of [3, Theorem 1.1]. Hence, $\partial_{\tau}C^{A}$ is continuous in $\Omega_{T_{c}}$, and consequently, $\partial_{\tau}C^{A}\geqslant 0$ everywhere in $\Omega_{T_{c}}$.
The same argument applies to $P^{A}$. By following the same procedure, it can be obtained that $\partial_{\tau}P^{A}\geqslant 0$ almost everywhere in $\Omega_{T_{p}}$. Theorem 1.1 in [3] yields that $\partial_{\tau}P^{A}$ is continuous, and hence nonnegative everywhere in $\Omega_{T_{p}}$.
Proof of Part (iv)
@@
D. Conditions of the operator $\mathcal{F}$ in Theorem 3.1
Let us prove that the operator $\mathcal{F}:\mathcal{A}\to\mathcal{X}$ in the proof of Theorem 3.1 satisfies
@@
    $\mathcal{F}(\mathcal{A})\subset\mathcal{A}$,
@@
    $\mathcal{F}(\mathcal{A})$ is precompact in $\mathcal{X}$.
(1) Fix any $w\in\mathcal{A}$ and consider $u=\mathcal{F}(w).$ Note that $J_{\varepsilon}(\tau,x)\geqslant 0,$ $\beta_{\varepsilon}(\cdots)\geqslant 0$ and $\nu\cdot D_{x}u(\tau,x)\leqslant 0$ for each $(\tau,x)$ at the lateral boundary and the inward pointing normal vector $\nu$ at $(\tau,x).$ Then by the comparison principle [17, 13p], we have $u\geqslant 0.$
@@
$\text{Let}\ \{w_{j}\}\$$\text{be a sequence in}\ \mathcal{A}\ \ \text{such that}\ \displaystyle{\lim_{j\to\infty}w_{j}}=w.$
@@
Then the $W^{1,2}_{p}$-estimate combined with the embedding theorem implies that
$\|u-u_{j}\|_{C(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}\|\beta_{\varepsilon}^{\prime}(\cdots)(w-w_{j})\|_{L^{\infty}(\Omega_{T}^{n})}$
for some constant $\mathcal{K}>0$. Therefore, it follows that ${\lim_{j\to\infty}u_{j}}=u.$
(3) Let $\{u_{j}\}_{j=1}^{\infty}$ be a sequence in the closure of $\mathcal{F}(\mathcal{A})$ with respect to the $\|\cdot\|_{C(\overline{\Omega_{T}^{n}})}$ norm.
Case 1. For each $u_{j}\in\mathcal{F}(\mathcal{A})$, there exists a sequence of functions $\{w_{j}\}_{j=1}^{\infty}$ such that $\mathcal{F}(w_{j})=u_{j}$. By applying the $W^{1,2}_{p}$ estimate on $u_{j}$, we obtain
$\|u_{j}\|_{W^{1,2}_{p}(\Omega^{n}_{T})}$ $\leqslant\mathcal{K}\big(\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(w_{j}-J_{\varepsilon})\|_{L^{p}(\Omega^{n}_{T})}+(e^{n}+e^{-n})T\big)$
$\leqslant\mathcal{K}\big(\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(-J_{\varepsilon})\|_{L^{p}(\Omega^{n}_{T})}+(e^{n}+e^{-n})T\big)$
@@
Case 2. For each fixed $u_{i}\notin\mathcal{F}(\mathcal{A})$, there exists a sequence $\{u^{(i)}_{l}\}_{l=1}^{\infty}\subset\mathcal{F}(\mathcal{A})$ such that
@@
Hence, there exists $\{w^{(i)}_{l}\}_{l=1}^{\infty}$ in $\mathcal{A}$ such that $\mathcal{F}(w^{(i)}_{l})=u^{(i)}_{l}$. Again, by the $W^{1,2}_{p}$ estimate, the embedding theorem, and the monotonicity of $\beta_{\varepsilon}$, we deduce $\|u_{l}^{(i)}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$ for some $\alpha\in(0,1)$. Take $l\rightarrow\infty$ to get $\|u_{i}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega_{T}^{n}})}\leqslant\mathcal{K}^{\prime}$.
@@
for all $j\in\mathbb{N}$.
By the Arzela-Ascoli theorem, there exists a subsequence $\{u_{j_{k}}\}_{k=1}^{\infty}\subset\{u_{j}\}_{j=1}^{\infty}$ and $u\in C(\overline{\Omega_{T}^{n}})$ such that
@@
Observe that $u$ is contained in the closure of $\mathcal{F}(\mathcal{A})$ by the assumption. This proves that $\mathcal{F}(\mathcal{A})$ is precompact in $\mathcal{X}=C(\overline{\Omega_{T}^{n}}).$